OpenStudy (anonymous):

Convert the angle from degrees to radians. a. 60° and b. -135°

OpenStudy (anonymous):

To go from degrees to radians you multiply it by pi and divide by 180 degrees. You can check this by looking at a circle, using previous knowledge of the unit circle. One full rotation is 2 pi, or 360 degrees. 2 pi/360 is pi/180. Also, you could check by inserting the degrees and radian equivalent into a trig. function on a calculator to see if you get the same answer (set the correct mode, radians or degrees).

OpenStudy (♪chibiterasu):

\[radians = \frac{ degrees \times \pi }{ 180 }\]

OpenStudy (anonymous):

(degree*pi)/180 soo \[\frac{ 60\pi }{ 180 }\]

OpenStudy (anonymous):

lolol jk just realized other people commented as I was trying to type that out xD.....carry on.

OpenStudy (anonymous):

hahah thanks y'all! I got it now :D